# MoLAN: A Unified Modality-Aware Noise Dynamic Editing Framework for Multimodal Sentiment Analysis

### v1

@@ Appendix C Implementation Detail
All experiments are conducted using the PyTorch framework on a hardware setup with 8 RTX A6000 GPUs. For the framework experiments, we strictly follow the hyperparameter settings reported in the original papers to ensure fair and consistent comparisons. In the methodological experiments, all models are trained using the AdamW optimizer. Specifically, the learning rate is set to 5e-6 for the MOSI and MOSEI datasets, 1e-5 for the CH-SIMS dataset, and 2e-8 for the IEMOCAP dataset. To ensure the stability of the experimental results, we fix the random seed to 1 and run each experiment five times independently, reporting the average result.
Appendix D More Ablation Experiments
MoLAN Framework
As shown in Table 9, we conduct ablation experiments on different parts of the MoLAN framework, including: (1) Using denoising methods only on video or only on audio modality. (2) Using 1D blocks for both video and audio. (3) Using 2D blocks for both video and audio. (4) Using 1D blocks for video and 2D blocks for audio.
Experimental results show that simultaneous block processing of visual and audio information can effectively filter out noise interference in each modality, thereby obtaining a more robust and discriminative feature representation. In further comparative experiments, we use multiple block dimensions. When the visual features are divided into two dimensions and the audio features are divided into one dimension, the model performance is optimal. This is highly consistent with the inherent structural characteristics of the two types of modalities. On the contrary, if the visual or audio information is divided in a block manner that does not match its information characteristics, it will lead to incomplete information after segmentation and loss of local features, which will lead to performance degradation. It can be seen that our block strategy improves denoising effects.
Table 9: MoLAN Framework Experiments.Take MulT as an example. 1D, 2D denote one-dimensional blocking and two-dimensional blocking, respectively.
Denoise-Driven Contrastive Learning
@@ Appendix D More Ablation Experiments
To further verify the effectiveness of denoising-driven contrastive learning (DC) in modality denoising, we use two metrics in the field of contrastive learning: alignment and uniformity. Alignment measures the average distance of the same positive sample pair in the embedding space. The smaller the value, the more closely the sample features are clustered. Uniformity reflects the degree of dispersion of all samples on the unit hypersphere. The lower the value, the greater the distance between negative sample pairs, and the noise is easier to distinguish.
@@
Table 10: Ablation studies of Denoise-Driven Contrastive Learning.
As shown in Table 10, DC can enhance modality discrimination. After introducing DC, the alignment of positive sample pairs is further reduced, indicating that homologous samples are more closely clustered in the feature space. At the same time, the uniformity of negative sample pairs is significantly improved, indicating that the model’s ability to distinguish noisy samples has been greatly improved. This shows that DC plays a key role in the denoising process, enabling the model to learn denoising feature representations more efficiently, thereby improving the performance of MSA task.

### v2

@@ B.3 Implementation Detail
All experiments are conducted using the PyTorch framework on a hardware setup with 8 RTX A6000 GPUs. For the framework experiments, we strictly follow the hyperparameter settings reported in the original papers to ensure fair and consistent comparisons. In our methodological experiments, to ensure fair comparison, we maintain the same parameter size settings for our modality encoders as in previous studiesWu et al. (2024). Model is trained using the AdamW optimizer to achieve stable and efficient convergence performance. Specifically, the learning rate is set to 5e-6 for the MOSI and MOSEI datasets, 1e-5 for the CH-SIMS dataset, and 2e-8 for the IEMOCAP dataset. In the framework experiments, we first fine-tune the MLLM using LoRAHu et al. (2022a) on MSA datasets to adapt it to the task characteristics and stabilize model performance. We then integrate the fine-tuned model into the proposed framework to verify its effectiveness. In the method experiments, the MLLM results report the performance of the model after LoRA fine-tuning. To ensure the stability of the experimental results, we fix the random seed to 1 and run each experiment five times independently, reporting the average result.
Appendix C Significance Testing
For Table 2, we perform significance testing on seven experimental groups, with p-value of $4.26\times 10^{-2}$, $6.01\times 10^{-4}$, $2.95\times 10^{-3}$, $08\times 10^{-3}, 1.1$$19\times 10^{-2}$, $4.28\times 10^{-3}$, and $4.45\times 10^{-3}$, all of which << 0.05 indicate significant differences.
For Table 4, the significance testing on MMML and MoLAN+, with p-value (0.01 << 0.05) indicating significant differences.
Appendix D More Experiments
Table 8: Ablation study of denoising strength computation guidance.
D.1 Block Partitioning Strategy Selection
We adopt a block partitioning strategy based on the factor closest to the square root of the feature dimension, since not all feature dimensions are perfect squares. This design ensures a more balanced division of feature regions and avoids the two extremes of excessively large or excessively small blocks. Specifically, when the block is too large, it tends to mix effective information with noise within the same region, making it difficult for the denoising strength to accurately distinguish between the two. This may result in incomplete noise removal or excessive suppression of critical information. In contrast, when the block is too small, it over-segments semantic structures, weakens inter-block correlations. By comparison, the close to square root block partitioning strategy achieves a better trade-off between signal–noise separation and semantic integrity, allowing the denoising process to retain critical information while removing noise.
As shown in Table 9, we design comparative experiments with different block sizes for both visual and audio modalities. The experimental results show that, in both visual and audio feature processing, excessively large or small block partitioning lead to a decline in model performance. In contrast, using the close to square root block partitioning strategy achieves the best balance between critical information preservation and noise suppression, allowing the model to capture critical features more effectively and maintain overall performance stability. Overall, the proposed close to square root block partitioning strategy aligns well with the characteristics of each modality and lays a solid foundation for subsequent denoising.
Figure 4: Performance comparison under different similarity thresholds $\theta$. The four curves represent experimental results on CMU-MOSI, CMU-MOSEI, CH-SIMS, and IEMOCAP datasets, respectively. The x-axis denotes the similarity threshold $\theta$, and the y-axis indicates the model performance.
D.2 Similarity Threshold Analysis
To determine the optimal similarity threshold $\theta$ in the noise-suppressed cross attention mechanism, we conduct a series of comparative experiments across four benchmark datasets, as shown in Figure 4. The threshold $\theta\in[0,1]$ controls the activation of the denoising mask by determining which sub-blocks are preserved or suppressed based on similarity scores.
The experimental results demonstrate a consistent trend across all datasets: as $\theta$ increases from 0.3 to 0.8, model performance initially improves and then declines. A smaller $\theta$ (e.g., 0.3–0.4) allows excessive noisy features to pass through, resulting in incomplete denoising and suboptimal performance. Conversely, an overly large $\theta$ (e.g., 0.7–0.8) overly suppresses critical information, causing semantic loss and degraded performance. The performance peaks at ($\theta$ = 0.5), where the model achieves the best trade-off between noise suppression and information preservation. Therefore, we set $\theta=0.5$ as the default similarity threshold in all experiments.
Table 9: Block Sizes Experiments.Take MulT+MoLAN on the CMU-MOSI and CMU-MOSEI dataset as an example.The audio feature dimension is $[128,375,20]$, and the visual feature dimension is $[128,500,20]$.The number after audio indicates the one-dimensional block size $(j)$, and the number after video indicates the two-dimensional block size $(k,j)$.
D.3 Denoising-Driven Contrastive Learning Analysis
@@
To further verify the effectiveness of denoising-driven contrastive learning (DC) in modality denoising, we use two metrics in the field of contrastive learning: alignment and uniformityWang and Isola (2020). Alignment measures the average distance between positive sample pairs in the embedding space, reflecting the model’s ability to cluster semantically similar samples. A smaller value indicates tighter intra-class compactness. Uniformity measures how evenly all samples are distributed on the unit hypersphere. A smaller value implies larger inter-class separation and better discrimination against noise.
As shown in Table  10, after incorporating DC, the model achieves lower alignment on both CMU-MOSI and CMU-MOSEI datasets, suggesting that positive samples cluster more tightly in the embedding space. Meanwhile, the uniformity score decreases significantly, indicating that negative samples become more separable and noise becomes easier to distinguish. These results indicate that DC effectively improves the model’s ability to distinguish critical information features from noisy redundant features, enabling the MSA model to learn purer modality representations.
@@
Table 10: Ablation studies of Denoising-Driven Contrastive Learning. Alignment and Uniformity metrics on CMU-MOSI and CMU-MOSEI datasets.
D.4 Denoising strength computation guidance
In our design, the denoising strength of each block is determined based on the text modality. To validate this choice, we conduct an ablation study comparing different guiding modalities, as shown in Table 8. The results show that using text as the guidance yields the best performance on two datasets. In contrast, using visual or audio features as the guidance leads to a performance drop, indicating that these modalities introduce more noise. These findings are consistent with the results observed in pilot study (Table 7), further confirming that text-guided denoising offers more reliable representations for MSA.
Figure 5: Case Study. The blue area in the visual modality represents the target person. The red, green, and blue colors in the audio modality represent the attention distribution of MMML, t-HNE, and MoLAN+ on different audio segments, respectively. The red boxes in the audio mark the noisy segments. The heatmap shows the model’s attention strength in different visual regions.
Appendix E Case Study
As shown in Figure 5, we present two representative case studies comparing MoLAN+ with the original SOTA model MMML and the latest denoising model t-HNE. In the left, we compare MoLAN+ with MMML. It can be observed that MMML pays almost equal attention to different visual and audio segments without effectively suppressing noise interference. Consequently, the model fails to distinguish the emotionally relevant regions and misinterprets the sentiment. In contrast, MoLAN+ focuses more accurately on the target speaker and the emotionally relevant regions, leading to a correct sentiment analysis. In the right, we compare MoLAN+ with t-HNE. Although t-HNE performs denoising, it applies a uniform denoising intensity across all modalities, which results in the loss of critical information necessary for MSA. This over-smoothing effect causes the model to generate incorrect sentiment. Conversely, MoLAN+ adaptively balances noise suppression and information retention through its dynamic noise-editing mechanism, thereby capturing emotion-relevant cues and achieving a more reliable analysis.